There is 1.5 L of water in a kettle on a cooking plate of 2.0 KW. It takes 7 minutes to heat the water from 20 C to the poiling point.
a) How much electrical energy is required ?
b) What is the efficiency of the heating process

Answer by Jk22(389)

a) 14/60 kWh = 0.84 MJ
b) the heat capacity of water is 4180 J/kg/K, hence 1.5 l that is heated by 80 degrees needs an energy of : 4180*1.5*80 = 0.5016 MJ

hence the efficiency of the process is : (0.84 - 0.5016)/0.5016 = 67 %